OpenStudy (anonymous):

help, please 4x + 5y + z = $23 6x + 6y = $30 3x + 7y + 5z = $28

OpenStudy (anonymous):

you looking for x & y on all of these ?

OpenStudy (anonymous):

and z

OpenStudy (anonymous):

from the second : (let me ignore the $ sign) x + y = 5 so x = 5-y plug it into the first: 4(5-y) +5y + z =23 20 - 4y + 5y +z = 23 y + z = 3 y = 3-z so we have y= 3-z so we can write x = 5-y in terms of z : x = 5-(3-z) = 2 + z plug it into the last equation: 3(2+z) +7(3-z) + 5z = 28 27 +z = 28 z = 1$ so x= 3$ and y= 2$